Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, November 10, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 10, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/10/98 (Tuesday)

Segment 1 (Emidio to ANF) Report for 11/10/98. An excavator dug on the north side of the large culverts at Peace Valley road ~EMP 25.45. Work began at the skip at EMP 21.75, by clearing the ROW of vegetation and some grading. ON the Tejon Bench, A tie-in section of pipe was placed in trench at ~EMP 18.42. Once the tie-in is completed, that section will be backfilled and access to the rest of the ROW may be easier. Silt fence was repaired and straw bales placed along a section from EMP 18.16 to EMP 18.23. Approximately 100' of trench was opened near EMP 16.98. A section of pipe was welded together at EMP 16.95 for a trench segment immediately south of its location. Padding and Backfill continued along Access Road B. Fiber Optic Conduit was put in from EMP 7.07 to the west to ~EMP 6.6. Trenching and shoring along HWY 166 continued near EMP .05, while coating of the pipe sections here took place. Trenching on Grapevine road continued. Grapevine and Emidio station and electrical work continued, with Welding at Grapevine station. Bending operations took place at ~ EMP 14.65.

Segment 2 (ANF to Sylmar) Report for 11/10/98): 2C (Templin Hwy to Castaic): Work on the tie-in of the testing cutout at Templin Hwy included lowering-in, welding, and x-ray. The construction of the block wall on the ANF access road south of Templin Hwy has been completed. South of the I-5 bore, Revegetation occurred between 53.51 and 55.46, and included hydro seeding, curlex-laying, and placing C-350 matting in gully crossings. Coating reached EMP 55.46, and is now complete for Segment 2C except for the Templin cutout. Shading moved north to EMP 55.67, and backfill and F/O conduit-laying to 55.79. Work at the Lake Hughes block valve continued until 5:00 PM. 2B (Castaic to Sylmar): F/O cable continued to be pulled from Lake Hughes Parkway to Haslely Cyn. Boring continued at the Feedmill Road directional drill. At Calgrove Ave, work involved backfill and removal of shoring boxes. Potholing occurred south of Weldon Cyn at EMP 72.47. At Sierra Hwy, EMP 72.69, 150 feet of trench was excavated, and several cross lines exposed. Stringing, welding, and coating occurred south of Sierra Hwy near EMP 72.75. 220 feet of asphalt and 350 feet of F/O cable were laid, and slurry poured around the Mobile Oil line to prevent wall collapse, at EMP 73.34. Crews met with CalTrans personnel to discuss time protocol for the section south of Weldon cyn.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 10 November 1998: 3A: Minor activity occurred on this section today. Backfill worked in various areas on clean up. Work on marking the line continued, with installation of signs starting at Wolfskill in San Fernando. The crew at the Tuxford bore continued to prepare the area for the bore crew. 3B: Minor clean up. 3C: All welders from Segment 3A and 3C worked on the insertion of the 20" carrier pipe into the 36" casing pipe inside the aqueduct structure. This crew worked late. Backfill focused on the tie in of the FOC at Roxford Street, and closing up the south side. Also, slurry was poured around the pipe in San Fernando Road near the I-5 overpass, just south of Sepulveda Boulevard. Some clean up work occurred across the ROW. The bore at Hubbard resumed, with an additional 60' of casing pushed.

Segment 4B (El Segundo to Lynwood) 11-10-98: Backfilling activities continued at Santa Ana and Alameda bore pit ESEMP 0.04. Construction activities continued at El Segundo station ESEMP 11.00. This crew was pushing the pipe through the casing. The crew at Slater ESEMP 1.68 was installing a Hand Hole. The crew at Lynwood station EMP 0.0. was cutting pipes and installing electrical wires. The paving crew was working at 108 Th St. ESEMP 1.11. Fiberoptic testing occurred at 108 Th St ESEMP 1.14 and at 109 Th St. 1.20. The pipe gang at Alameda and Fernwood Av EMP 0.0 were trying to tie-in the pipe to the off-set that goes to the Lynwood station, however they were having great difficulties whit that tie-in. During backfilling the Western bore pit ESEMP 5.56 a section of the pipe got damaged, last week. This section of the pipe will be removed, and the backfilling process will be delayed for a week.

Segment 5 (Lynwood to Long Beach) Report for 10 November 1998: The directional drill at Highway 91/Compton Creek had a frac-out on the west bank of the creek. Drilling mud got into the creek before containment was complete. A memo was issued. Of the four bores at the Wilmington refinery, two are complete (one cased, one cased with pipe) and two are cased bores that are in progress.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page